# Fish Tracking, Counting, and Behaviour Analysis in Digital Aquaculture: A Comprehensive Review

### v1

@@ III-B Counting methods based on computer vision technology
This section summarizes and analyzes the current computer vision-based fish counting aquaculture methods, focusing on two main categories: image-based counting and video-based counting. The summary of the computer vision-based methods can be seen in Table. II. The subsections delve into the details of each category, discussing the advancements, challenges, and future directions in the field. By bridging the gap between laboratory-based experiments and real-world applications, computer vision-based fish counting can become an indispensable tool for sustainable aquaculture management.
III-B1 Image-based counting method
@@
TABLE II: Different counting methods based on computer vision.
@@
Amount Dataset Model Count points
@@
Tank 100 786 $\mathrm{MAN}$ Center Accuracy $97.12\%$
@@
Cage 62 200 $\mathrm{RCNN}$ Bounding Box Accuracy $92.4\%$
@@
Aquarium 9 -
Adaptive
Thresholding
Skeleton
Average
counting error
$6\%$
Solve the
overlapped-fish
problem cleverly
Only adapted to
relatively small
fish densities
[112]
Net Cage 250 1000 $\mathrm{PTV}$ Centroid Detection rate $90\%$
@@
Aquarium 100 600 LS-SVM Skeleton Accuracy $98.73\%$
@@
Aquarium 300 3200 MSENet Centroid MAE 3.33
@@
Early studies focused on detecting-based methods, which rely heavily on the accuracy of fish image segmentation from the background [117]. These methods, such as artificial neural networks (BPNN) [118], showed potential for automatic fish counting in scenarios with a limited number of fish. However, they often struggled with complex adhesions in fish images and overlapping fish [119, 111]. To address the challenges of overlapping fish, adaptive segmentation algorithms were developed to extract the geometric features of fish [114]. Combined with machine learning models like LS-SVM, these algorithms showed improved counting accuracy compared to BPNN models, particularly in scenarios with similar fish sizes and low stocking densities. However, the performance of these models declined when faced with high fish densities and changing geometric shapes due to fish overlap [120]. Further advancements in fish image segmentation were made by introducing more general adaptive thresholding methods and skeleton extraction-based methods to handle overlapping fish [112]. While these methods performed well under controlled laboratory conditions, their accuracy diminished in real-world aquaculture environments, where factors such as high fish school density, poor visibility, and insufficient light posed significant challenges [113].
@@
The introduction of deep learning techniques has opened new avenues for fish counting in aquaculture. With the increasing availability of fish datasets, deep learning models have been applied to this domain, offering strong adaptability and easy transformation without requiring complex feature extraction work [121, 107]. Convolutional neural networks (CNNs) have been shown to achieve high accuracy in detecting and counting fish of different sizes by adjusting different thresholds [109].
@@
Despite the advancements made in fish counting methods, several challenges remain. Density-based methods are sensitive to the degree of occlusion, with higher fish densities leading to greater errors. Moreover, variations in water quality, light conditions, camera angle, water depth, and surface refraction can cause significant differences in the appearance of fish across different farming environments, affecting the accuracy and generalization ability of counting models. To address these challenges, future research should create more comprehensive and diverse datasets that capture the variability encountered in real-world aquaculture settings. Efforts should also be directed towards improving counting accuracy, model generalization ability in high-density areas, and maintaining accuracy under different pond conditions.
III-B2 Video-based counting method
Video-based counting methods offer a more efficient alternative to counting objects in single images, enabling the development of reliable and inexpensive systems for counting fish in sequential videos [123]. However, directly applying current automatic detecting and counting frameworks in underwater environments presents several challenges. Firstly, underwater cameras are susceptible to contamination by impurities in the water, leading to deterioration of video quality. Additionally, current communication technology and cost limit underwater video real-time transmission technology, resulting in delays that impact real-time detection [124]. Secondly, underwater videos suffer from colour shift and contrast degradation due to light absorption and scattering in the water, making object detection and segmentation more difficult than in land-based applications [125].
To address these issues, numerous image enhancement algorithms have been proposed to improve the quality of underwater images [126]. These algorithms aim to restore the real colour and improve the contrast of underwater images. However, the effectiveness of these enhancement techniques varies greatly depending on the environment and lighting conditions [127]. Furthermore, the detection and segmentation results directly depend on the image enhancement and segmentation performance, which can be time-consuming.
Despite these challenges, video-based counting methods find applications in various aquatic environments, such as aquaculture and fisheries management. In aquaculture, these methods can be used to estimate fish catch and abundance statistics, reducing the time and effort required for manual recording by fishermen [128, 129]. In the context of stream fish research and management, video technology provides a new strategy for estimating fish abundance, although its effectiveness may vary depending on the age and behaviour of the fish species being studied [130, 131].
As technology advances and more robust algorithms are developed, video-based counting methods are expected to play an increasingly important role in accurately assessing fish populations in various aquatic environments. Combining video methods with machine learning models promotes powerful new directions for river fish research, management, and protection. Future research should focus on improving the reliability and efficiency of these methods while addressing the specific challenges posed by underwater video acquisition and analysis, such as image quality degradation, real-time transmission limitations, and the need for effective image enhancement techniques. By tackling these issues, video-based counting methods can provide valuable insights into fish populations and support sustainable fisheries management practices.
III-C Counting methods based on acoustic technology
@@ III-C1 Acoustic imaging methods
Multi-beam imaging sonar such as Adaptive Resolution Imaging Sonar (ARIS) and Dual-frequency Identification Sonar (DIDSON) are normally used to monitor migratory fish in rivers [132]. These systems produce high-resolution underwater sonar video output without the need for underwater light, allowing for fish counting and measuring directly from the footage, even in turbid waters and overnight [133].
DIDSON, developed by the Applied Physics Laboratory at the University of Washington [134], is a multi-beam sonar system frequently used to acquire underwater acoustic images for fish identification and counting. As DIDSON uses sound instead of light, it is not affected by water turbidity and can collect data during both day and night [135, 136]. However, studies have shown that manual counting of DIDSON data can be time-consuming and prone to errors, with large deviations between operators [137, 138]. This may be because Echoview repeatedly calculated nearly stationary horizontal positions within the DIDSON field of view [139].
Figure 6: Example of DIDSON image counting [140].
@@
The new generation of acoustic cameras includes Adaptive Resolution Imaging Sonar (ARIS) (Sound Metrics Corp, WA, USA), which operates at higher frequencies compared to DIDSON, offering greater flexibility and improved image resolution [149, 150]. A comparison of fish monitoring data based on the ARIS sonar system and the GoPro camera showed that the detection rate of the sonar-based system was 62.6% (compared to the amount captured by the net), exceeding the 45.4% of the camera-based system [151].
@@
Table III summarizes the advantages and disadvantages of sonar imaging counting methods in aquaculture and their practical applications. Despite the limitations, acoustic counting methods remain valuable for monitoring fish populations in challenging underwater environments where visual counting methods may be impractical or ineffective.
TABLE III: A comparison of different methods based on acoustic imaging.
@@ IV Fish school behaviour analysis
Fish behaviour, a direct result of the living environment and growth state, includes both normal (i.e. feeding behaviour, swimming behaviour, reproduction behaviour, gathering behaviour) and abnormal behaviours (i.e. disease behaviour, hypoxia behaviour, cannibalism behaviour) [167, 168, 169]. Poor water quality and management in aquaculture can cause fish stress behaviour [170], leading to immune suppression, slow growth, and reduced productivity and welfare [171]. Traditional fish behaviour analysis, relying on human observers, is often unreliable, time-consuming, and labour-intensive [5, 6]. Accurate estimation of fish behaviour is crucial for optimizing resource use, controlling water quality, and improving fish welfare and economic benefits [172]. The following sections will explore the latest advancements in fish behaviour analysis, focusing on computer vision-based methods for assessing fish school behaviour and feeding behaviour, providing insights into the current state of the art and potential future directions for research and application in this field.
@@ IV-A1 Fish feeding behavior
Existing approaches mainly use digital cameras to capture the corresponding images as input and characterize the fish behaviour with discrete feeding intensity (e.g., “None”, “Weak”, “Medium” and “Strong[178, 179, 180, 181]) as a classification problem, modelled by Convolutional Neural Networks (CNNs). However, fish-feeding behaviour is a dynamic and continuous process. Single images are insufficient to capture the context of fish feeding intensity [177, 182]. As an alternative, video-based methods have been proposed to exploit spatial and temporal visual information for FFIA, which offers rich context for capturing fish feeding behaviour. Raw RGB videos were converted into optical flow image sequences and fed into a 3D convolutional neural network (3D CNN) to evaluate fish feeding intensity, achieving a very high accuracy [183, 184].
@@ IV-A3 Other abnormal behavior
While computer vision and image processing technologies offer economical and effective means for monitoring abnormal fish behaviour, the relative scarcity of abnormal behaviour data has hindered in-depth research. Most existing studies have been conducted in controlled laboratory environments, which may not accurately represent the complex factors in real-world aquaculture settings [200, 201]. Overcoming the challenges posed by complex water environments, uneven lighting, large numbers of individuals, and intricate fish movements is crucial for developing robust and reliable multi-target abnormal behaviour monitoring and tracking systems in computer vision [61].
IV-B Fish school behaviour analysis based on Trajectory analysis
@@
Visual-based monitoring systems for detecting abnormal fish behaviour often rely on known scenes and predefined movement models, which can be subjective and lack adaptability to different environments [202]. Analyzing many target trajectories in a specific scene can reveal behaviour patterns and construct effective motion behaviour models with greater universality and applicability [182].
@@
Using visual monitoring systems, researchers can obtain 3D time-varying trajectory data (location and time information) of fish. Studies have tracked zebrafish using YOLOV2 and Kalman filters, obtaining movement trajectories that showed significantly faster swimming, greater agitation, and agglomeration in the centre of the aquarium during feeding periods [29]. Other researchers have developed semi-automatic in situ tracking systems to reconstruct synchronized 3D movement trajectories of individual reef fish in social groups, analyzing their behaviour when capturing plankton prey [37].
@@
However, relatively few studies on abnormal fish behaviour use trajectory analysis in aquaculture. This scarcity can be attributed to the limited availability of open datasets on abnormal fish behaviour and the rare occurrence of such behaviours, making data acquisition challenging. Moreover, fish trajectories inherently contain information about position, speed, and direction, and the definition of an abnormal trajectory may encompass multiple aspects. In the past decade, track-based anomaly detection methods have primarily relied on traditional clustering methods or focused on statistical models of trajectories. In contrast, the representation of trajectories remains an open problem [203, 204].
@@
To address these challenges and advance aquaculture, future research should draw inspiration from the successful application of abnormal trajectory behaviour analysis methods in other fields, such as crowd and vehicle monitoring. By adapting these techniques to fish’s unique characteristics and environments, researchers can create more powerful and flexible models for identifying and comprehending abnormal fish behaviour in various aquaculture contexts.

### v2

@@ III-B Fish counting methods based on computer vision technology
The current computer vision-based fish counting aquaculture methods focus on two main categories: image-based counting and video-based counting, as shown in Table. III. By bridging the gap between laboratory-based experiments and real-world applications, computer vision-based fish counting can become a useful tool for sustainable aquaculture management. The subsections provide more details about each category, the advancements, challenges, and future directions in the field.
III-B1 Image-based fish counting method
@@
TABLE III: Different counting methods based on computer vision.
@@
Fish
amount
Dataset Model Count points
@@
Tank 100 786 $\mathrm{MAN}$1 Center Accuracy $97.12\%$
@@
Cage 62 200 $\mathrm{RCNN}$4 Bounding Box Accuracy $92.4\%$
@@
Net Cage 250 1000 $\mathrm{PTV}$5 Centroid Detection rate $90\%$
@@
Aquarium 100 600 LS-SVM6 Skeleton Accuracy $98.73\%$
@@
Aquarium 300 3200 MSENet7 Centroid MAE8 3.33
@@
  • 1
    MAN (Multi-modules and attention mechanism)
  • 2
    DG-LR (Image density grading and local regression)
  • 3
    R2 (Coefficient of determination)
  • 4
    RCNN (Region-based convolutional neural network)
  • 5
    PTV (Particle tracking velocimetry)
  • 6
    LS-SVM (Least squares support vector machine)
  • 7
    MSENet (A lightweight network based on SENet )
  • 8
    MAE (Mean absolute error )
Early studies focused on detecting-based methods, which rely heavily on the accuracy of fish image segmentation from the background [161]. These methods, such as artificial neural networks (BPNN) [162], showed potential for automatic fish counting in scenarios with a limited number of fish. However, they often struggled with complex adhesions in fish images and overlapping fish [163, 156]. To address the challenges of overlapping fish, adaptive segmentation algorithms were developed to extract the geometric features of fish [158]. Combined with machine learning models like LS-SVM, these algorithms showed improved counting accuracy compared to BPNN models, particularly in scenarios with similar fish sizes and low stocking densities. However, the performance of these models degrades when faced with high fish densities and changing geometric shapes due to fish overlap [164]. Further advancements in fish image segmentation were made by introducing more general adaptive thresholding methods and skeleton extraction-based methods to handle overlapping fish [165]. While these methods performed well under controlled laboratory conditions, their accuracy diminished in real-world aquaculture environments, where factors such as high fish school density, poor visibility, and insufficient light posed significant challenges [157].
@@
Introducing deep learning techniques has opened new avenues for fish counting in aquaculture. With the increasing availability of fish datasets, deep learning models have been applied to this domain, offering strong adaptability and easy transformation without requiring complex feature extraction work [166, 152]. Convolutional neural networks (CNNs) have achieved high accuracy in detecting and counting fish of different sizes by adjusting different thresholds [154].
@@
Despite the advancements made in fish counting methods, several challenges remain. Density-based methods are sensitive to the degree of occlusion, with higher fish densities leading to greater errors. Moreover, variations in water quality, light conditions, camera angle, water depth, and surface refraction can cause significant differences in the appearance of fish across different farming environments, affecting the accuracy and generalization ability of the counting models. To address these challenges, future research should create more comprehensive and diverse datasets that capture the variability encountered in real-world aquaculture settings. Efforts should also be directed towards improving counting accuracy, model generalization ability in high-density areas, and maintaining accuracy under different pond conditions.
III-B2 Video-based fish counting method
Video-based counting methods offer a more comprehensive approach to enumerating fish than static image-based techniques, leveraging temporal information and movement patterns [168]. These methods can be broadly categorized into frame-by-frame analysis, tracking-based counting, and segmentation-based counting [9].
The frame-by-frame analysis based methods extend the image-based techniques to video sequences, applying detection and counting algorithms to each frame. Zhao et al. [33] proposed a method using background subtraction and object detection to count fish in video streams, achieving higher accuracy than single-image based counting methods due to the ability to average counts over multiple frames. Like this work, Pai et al. [169] used the YOLOv5 model [60] to detect and count fish in video streams and get similar results. This approach is relatively simple to implement and can leverage existing image-based algorithms. However, it does not fully use temporal information, may struggle with rapid fish movements between frames, and is prone to double-counting or missing fish due to occlusions.
Tracking-based counting methods follow individual fish across frames, offering insights into fish movement patterns and behaviours. In addition, the tracking-based counting methods can deal with the temporary occlusions, and reduce double-counting errors [30]. Albuquerque et al. [153] and Zhou et al. [170] developed a system combining object detection with multiple object tracking to count fish in dynamic videos for large-scale counting in practical production. This method uses segmentation and association, and can effectively solve the problem of mis-detection caused by adhesive fingerlings in mechanical counters (as shown in Fig. 8). However, in practical applications, these algorithms often struggle to achieve a balance between counting accuracy and processing speed, limiting their applicability in real-time commercial settings. Addressing this challenge, Zhang et al. [160] proposed a novel combination of the SORT algorithm [33] and YOLOv5-Nano [171] for fry tracking and counting. This innovative approach achieves accurate dynamic counting with even faster running speeds, representing a significant step towards real-time fish counting in aquaculture environments. Integrating lightweight deep learning models with efficient tracking algorithms showcases a promising direction for future research, potentially enabling more widespread adoption of video-based counting methods in commercial aquaculture operations.
Figure 8: Overview of the fry counting system. Fish is poured into a channel with water flow, LED lights provide stable illumination. An industrial camera records the fry swimming process and transmits it to a computer (or edge device), and the device counts the number of fry passed in real-time through the designed method.
Segmentation-based counting methods focus on isolating fish from the background in each frame and then counting the segmented objects. Instance segmentation methods have emerged as a promising solution for effectively segmenting individual fish in dense scenes, and combining these methods with advanced detectors has the potential to achieve real-time performance [172]. Wang et al. [173] propose a lightweight instance segmentation model based on YOLOv8 [63], realizing high-density and multi-target fish segmentation and counting in an occluded environment. This approach demonstrates the potential of segmentation-based methods in handling complex aquaculture scenarios. However, the segmentation performance is notably diminished in scenarios of significant occlusions, presenting a significant challenge for dense fish populations. To address this limitation, recent advancements in computer vision, such as the Segment Anything Model (SAM) [174], offer new possibilities. The SAM model is pre-trained on a large-scale dataset and exhibits exceptional robustness in segmentation tasks across various domains. Although no literature is found for fish counting based on SAM, its excellent segmentation capabilities hold promise for future applications in aquaculture. Looking ahead, integrating the robust segmentation abilities of SAM with specialized fish counting algorithms could potentially overcome the limitations of current methods, particularly in highly occluded environments. It could be a promising direction to adapt SAM for fish-specific segmentation tasks and optimise its real-time performance to meet the demands of commercial aquaculture operations.
III-C Fish counting methods based on acoustic technology
@@ III-C1 Acoustic imaging methods
Multi-beam imaging sonar such as Dual-frequency Identification Sonar (DIDSON) and Adaptive Resolution Imaging Sonar (ARIS) are normally used to monitor migratory fish in rivers [175]. These systems produce high-resolution underwater sonar video output without the need for underwater light, allowing for fish counting and measuring directly from the footage, even in turbid waters and overnight [176].
DIDSON [177] is a multi-beam sonar system frequently used to acquire underwater acoustic images for fish identification and counting (as shown in Fig. 9). As DIDSON uses sound instead of light, it is not affected by water turbidity and can collect data during both day and night [178, 179]. However, studies have shown that manual counting of DIDSON data can be time-consuming and prone to errors, with large deviations between operators [180, 181]. This may be because Echoview repeatedly calculated at nearly stationary horizontal positions within the DIDSON field of view [182].
Figure 9: Example of DIDSON image counting (This figure was reproduced from [183] with minor changes).
@@
The new generation of acoustic cameras includes ARIS (Sound Metrics Corp, WA, USA), which operates at higher frequencies compared to DIDSON, offering greater flexibility and improved image resolution [192, 193]. A comparison of fish monitoring data based on the ARIS sonar system and the GoPro camera showed that the detection rate of the sonar-based system was 62.6% (compared to the amount captured by the net), exceeding the 45.4% of the camera-based system [194].
@@
Table IV summarizes the advantages and disadvantages of sonar imaging counting methods in aquaculture and their practical applications. Despite the limitations, acoustic counting methods remain valuable for monitoring fish populations in challenging underwater environments where visual counting methods may be impractical or ineffective.
TABLE IV: A comparison of different methods based on acoustic imaging.
@@ III-C2 Hydroacoustic methods
III-D Comparative analysis and trends in fish counting methods
Fish counting methods in aquaculture can be broadly categorized into sensor-based, computer vision-based, and acoustic-based approaches, each with distinct advantages and limitations. Sensor-based counting, including infrared and resistivity counters, offers real-time data but may be limited by water conditions and fish density [9]. Infrared counters are non-invasive but struggle in turbid waters, while resistivity counters can operate in various conditions but may require fish to pass through specific channels [132, 13]. Computer vision-based counting is divided into image-based and video-based methods. Image-based counting is cost-effective and can handle large datasets but may struggle with overlapping fish [153]. Video-based counting offers more dynamic analysis, and better handling of movement and overlaps, but requires more computational power [169, 170]. Acoustic-based counting excels in challenging environments, including acoustic imaging and hydroacoustic methods [3]. Acoustic imaging (e.g., DIDSON) works well in turbid waters and at night but can be costly [179, 182]. Hydroacoustic methods are effective for quantifying fish density and biomass at large scales but may lack species-specific accuracy [199, 205].
Analysis of research literature reveals that the selection of fish tracking methods is closely tied to aquaculture environments and infrastructure types. In intensive pond and tank systems, computer vision-based methods have been widely studied and implemented [210], demonstrating particular success in controlled environments. Research in large-scale marine farming has explored various approaches, including acoustic methods for deep-water applications [197]. Recent studies have increasingly focused on AI-integrated tracking systems [22], while research in open-water aquaculture settings has predominantly investigated acoustic tracking methods due to their effectiveness in larger volumes [211]. The adoption trends are shaped by environmental conditions, species being counted, regulatory requirements, and economic factors. Due to cost constraints, developing regions often rely on manual counting or basic sensor technology. Welfare considerations are increasingly important in counting method selection. Noninvasive methods like computer vision and acoustic imaging are preferred to minimize fish stress. Ongoing research focuses on improving accuracy while reducing the impact on fish behaviour.
III-E Real-world applications in fish counting
Fish counting is a critical aspect of aquaculture management, essential for controlling production volume, feed usage, and regulatory reporting. Yanmar Marine Systems Co., Ltd. (YMS) developed a real-world application (as shown in Fig. 10) of automated fish counting to address the challenges faced in tuna farming  [^1] https://www.yanmar.com/global/about/technology/vision2/fish_counting_system.html. Traditionally, farm fish counting relied on labour-intensive and potentially inaccurate methods such as visual counting from camera footage or subtractive counting of dead fish. To overcome these limitations, YMS developed an automated fish counting system aimed at achieving 98% or higher accuracy in real-time tuna counting. This system integrates advanced image recognition technologies and consists of underwater cameras, a dedicated image processing computer, and specialized software. Key features include real-time counting using simultaneous imaging and count recognition, detection of environmental disturbances, and remote adjustment for optimal imaging conditions. The system is beneficial when introducing young Pacific bluefin tuna into nets, during the division of fish into multiple nets during growth phases, and for daily inventory management and feed usage monitoring. In testing, it achieved the target of 98% or higher accuracy, leading to its commercial release in December 2020. This automated system significantly reduces the workload on farm workers while maintaining high accuracy, demonstrating the practical application of computer vision and image processing techniques in commercial aquaculture settings and potentially revolutionizing farm management practices.
Figure 10: Example of the counting method developed by YMS.
Another typical example is a large commercial tilapia hatchery that implemented an advanced dynamic fry counting system using 32 cm wide channels (as shown in Fig. 8), high-resolution cameras, and a central server running YOLOv5-Nano [171] detection model and an improved SORT algorithm [160]. The system uses several innovative approaches to address common challenges such as occlusion, high density, and adhered fry. The improved SORT algorithm maintained consistent tracking of individual fry across frames, even in high-density situations. The detection model was trained on diverse datasets including edge cases to handle adhered fry. The optimized channel design, with careful calibration of width and inclination, encouraged fry separation and reduced occlusion. Temporal information from multiple frames was utilized to resolve ambiguities caused by occlusion or high density. After six months, the incubator reported significant improvements: fry count accuracy increased to within 3-4% (from 20% variance), counting time was reduced by 85%, overfeeding decreased by 12%, and fry survival rates improved by 3%. The implementation led to a 15% reduction in operational costs, a 10% increase in production efficiency, a 70% decrease in counting-related labour costs, and a 20% reduction in lost sales. Improved filtration and model fine-tuning addressed challenges introduced by turbid water and varying fry strains. This application demonstrates the potential of using advanced computer vision and tracking algorithms to enhance commercial aquaculture operations.
@@ IV Fish school behaviour analysis
Fish behaviour, a direct result of the living environment and growth state, includes both normal (i.e. feeding behaviour, swimming behaviour, reproduction behaviour, gathering behaviour) and abnormal behaviours (i.e. disease behaviour, hypoxia behaviour, cannibalism behaviour) [212, 213, 214]. Poor water quality and management in aquaculture can cause fish stress behaviour [215], leading to immune suppression, slow growth, and reduced productivity and welfare [216]. Traditional fish behaviour analysis, relying on human observers, is often unreliable, time-consuming, and labour-intensive [4, 5]. Accurate estimation of fish behaviour is crucial for optimizing resource use, controlling water quality, and improving fish welfare and economic benefits [217]. The following subsections will explore the latest advancements in fish behaviour analysis, providing insights into the current state of the art and potential future directions for research and application in this field.
@@ IV-A1 Fish feeding behavior
Existing approaches mainly use digital cameras to capture the corresponding images as input and characterize the fish behaviour with discrete feeding intensity (e.g., “None”, “Weak”, “Medium” and “Strong[223, 224, 225, 226]) as a classification problem, modelled by CNNs. However, fish-feeding behaviour is a dynamic and continuous process. Single images are insufficient to capture the context of fish feeding intensity [222, 227]. As an alternative, video-based methods have been proposed to exploit spatial and temporal visual information for fish feeding intensity assessment (FFIA), which offers rich context for capturing fish feeding behaviour. Raw RGB videos were converted into optical flow image sequences and fed into a 3D CNN to evaluate fish feeding intensity, achieving a very high accuracy [228, 229].
@@ IV-A3 Other abnormal behavior
While computer vision and image processing technologies offer economical and effective means for monitoring abnormal fish behaviour, the relative scarcity of abnormal behaviour data has hindered in-depth research. Most existing studies have been conducted in controlled laboratory environments, which may not accurately represent the complex factors in real-world aquaculture settings [245, 246]. Overcoming the challenges posed by complex water environments, uneven lighting, large numbers of individuals, and intricate fish movements are crucial for developing robust and reliable abnormal behaviour monitoring and tracking systems [78].
IV-B Fish behaviour analysis based on multi-object tracking
@@
Visual-based monitoring systems for detecting abnormal fish behaviour often rely on known scenes and predefined movement models, which can be subjective and lack adaptability to different environments [247]. Advanced tracking methods, however, offer more robust and adaptable solutions for analysing fish behaviour in diverse aquaculture settings [227].
@@
Tracking technologies enable researchers to obtain comprehensive data on fish movement, including average, maximum, and minimum speeds, acceleration, average collision frequency, and trajectory changes. For example, studies have tracked zebrafish using YOLOV2 [248] and Kalman filters, obtaining movement trajectories that showed significantly faster swimming, greater agitation, and agglomeration in the centre of the aquarium during feeding periods [29]. Similarly, semi-automatic in situ tracking systems have been developed to reconstruct synchronized 3D movement trajectories of individual reef fish in social groups, analysing their behaviour when capturing plankton prey [36].
Recent advancements in tracking algorithms have expanded the capabilities of fish behaviour analysis. Huang et al. [249] developed an early warning system for detecting nocardiosis in largemouth bass (Micropterus salmoides) based on YOLOv8 [63] and ByteTrack [67]. The system quantifies fish velocity and turning angles through trajectory analysis, enabling effective early disease prevention. Similarly, Zhao et al. [250] employed YOLOv8+ByteTrack multi-target tracking to monitor zebrafish behaviour under various environmental stressors, extracting key movement features (e.g. speeds, average collision frequency, and trajectory changes) to identify exposure to different pollutants rapidly. Xiao et al. [251] enhanced the tracking performance of largemouth bass (Micropterus salmoides) by combining an improved YOLOv8 [63] with ByteTrack, incorporating trajectory confidence information. This approach improved tracking accuracy and enabled more detailed analysis of fish swimming patterns. By examining the relationship between swimming speed, swimming time, and spatial coordinates, the researchers were able to quantify swimming ability in terms of both speed and endurance.
@@
Despite the advancements in tracking algorithms, their application to abnormal fish behaviour analysis in aquaculture remains limited. This scarcity can be attributed to several factors: limited availability of open datasets on abnormal fish behaviour, the rare occurrence of such behaviours making data acquisition challenging, the complexity of fish trajectories containing multi-dimensional information (position, speed, direction), and difficulty in defining ‘abnormal’ trajectories which may encompass multiple aspects[252, 253].
@@
To address these challenges and advance aquaculture, future research should focus on developing specialized datasets for abnormal fish behaviour in aquaculture settings, adapting advanced tracking algorithms to the unique characteristics of fish and aquatic environments, integrating domain knowledge from aquaculture experts to better define and detect abnormal behaviours, and drawing inspiration from successful applications of anomaly detection in other fields, such as crowd and vehicle monitoring. By focusing on these areas, researchers can create more powerful and flexible models for identifying and comprehending abnormal fish behaviour in various aquaculture contexts, leveraging the full potential of advanced tracking technologies.